OpenStudy (anonymous):

1. y = x + 6 y = 2x (1 point) (6, 12) (–12, –6) (–6, –12) (2, 4) 2. 3x + 2y = 7 –3x + 11 = y (1 point) (6, –3) (6, –7) (5, –4) 3. 2x – y = –7 4x – y = –4 (–1.5, 4) (1.5, 10) (4, –1.5) (–1.5, –2) For questions 4–5, how many solutions does the system have? 4. –4y + 4 = x 2x + 8y = 8 one solution two solutions infinitely many solutions no solution 5. 6x + 2 = y 3y – 18x = 12 one solution two solutions infinitely many solutions no solution

OpenStudy (anonymous):

For questions 1–3, what is the solution of the system? Use substitution. For questions 4–5, how many solutions does the system have?

OpenStudy (anonymous):

for 1-3, substitute y=2x into first equation. 2x=x+6 2x-x=6 x=6 substitute x=6 into 2nd equation y=2x y=2*6 =12 (6.12)